OpenStudy (blackstreet23):

Show 7that if a, b, c and d are integers, where! = 0, such that a | c and b | d, then ab | cd?

OpenStudy (solomonzelman):

My interpretation:\(\\[0.6em]\) \(\color{black}{{\rm Show{\small ~~}that{\small ~~}\forall}a,b,c,d\in\mathbb{Z},{\small~~}{\rm if}{\small~~}a|c{\small~~}{\rm and}{\small~~}b|d,{\small~~}{\rm then}{\small~~}ab|cd.}\)

OpenStudy (solomonzelman):

Proof: \(\\[1.4em]\) Since \(a|b\), therefore \(b=am\) for some \(m\in\mathbb{Z}\). Since \(c|d\), therefore \(d=cn \) for some \(n\in\mathbb{Z}\). Consequentially, \(cd=(am)(cn)=ac(mn)\), with \(mn\in\mathbb{Z}\). Therefore \(ab|cd\).